# Agentic Context Engineering: Evolving Contexts for Self-Improving Language Models

### v2

@@ 4.2 Baselines and Methods
Base LLM.
The base model is evaluated directly on each benchmark without any context engineering, using the default prompts provided by dataset authors. For AppWorld, we follow the official ReAct [yao2023react] implementation released by the benchmark authors, and build all other baselines and methods on top of this framework.
In-Context Learning (ICL) [agarwal2024many].
@@
MIPROv2 [opsahl2024optimizing].
MIPROv2 is a popular prompt optimizer for LLM applications that works by jointly optimizing system instructions and in-context demonstrations via bayesian optimization. We use the official DSPy implementation [DSPyMIPROv2], setting auto="heavy" to maximize optimization performance.
GEPA [agrawal2025gepa].
GEPA (Genetic-Pareto) is a sample-efficient prompt optimizer based on reflective prompt evolution. It collects execution traces (reasoning, tool calls, intermediate outputs) and applies natural-language reflection to diagnose errors, assign credit, and propose prompt updates. A genetic Pareto search maintains a frontier of high-performing prompts, mitigating local optima. Empirically, GEPA outperforms reinforcement learning methods such as GRPO and prompt optimizers like MIPROv2, achieving up to 10–20% higher accuracy with as much as 35× fewer rollouts. We use the official DSPy implementation [DSPyGEPA], setting auto="heavy" to maximize optimization performance.
Dynamic Cheatsheet (DC) [suzgun2025dynamic].
DC is a test-time learning approach that introduces an adaptive external memory of reusable strategies and code snippets. By continuously updating this memory with newly encountered inputs and outputs, DC enables models to accumulate knowledge and reuse it across tasks, often leading to substantial improvements over static prompting methods. A key advantage of DC is that it does not require ground-truth labels: the model can curate its own memory from its generations, making the method highly flexible and broadly applicable. We use the official implementation released by the authors [Suzgun2025_DynamicCheatsheet_code] and set it to use the cumulative mode (DC-CU).
ACE (ours).
ACE optimizes LLM contexts for both offline and online adaptation through an agentic context engineering framework. To ensure fairness, we use the same LLM for the Generator, Reflector, and Curator (non-thinking mode of DeepSeek-V3.1 [deepseekai2024deepseekv3technicalreport]), preventing knowledge transfer from a stronger Reflector or Curator to a weaker Generator. This isolates the benefit of context construction itself. We adopt a batch size of 1 (constructing a delta context from each sample). We set the maximum number of Reflector refinement rounds and the maximum number of epochs in offline adaptation to 5.
@@
Table 1: Results on the AppWorld Agent Benchmark. "GT labels" indicates whether ground-truth labels are available to the Reflector during adaptation. We evaluate the ACE framework against multiple baselines on top of the official ReAct implementation, both for offline and online context adaptation. ReAct + ACE outperforms selected baselines by an average of 10.6%, and could achieve good performance even without access to GT labels.
4.3 Results on Agent Benchmark
Analysis.
As shown in Table 1, ACE consistently improves over strong baselines on the AppWorld benchmark. In the offline setting, ReAct + ACE outperforms both ReAct + ICL and ReAct + GEPA by significant margins (12.3% and 11.9%, respectively), demonstrating that structured, evolving, and detailed contexts enable more effective agent learning than fixed demonstrations or single optimized instruction prompts. These gains extend to the online setting, where ACE continues to outperform prior adaptive methods such as Dynamic Cheatsheet by an average of 7.6%.
@@ 4.3 Results on Agent Benchmark
Notably, on the latest AppWorld leaderboard (as of September 20, 2025; Figure 5), on average, ReAct + ACE (59.4%) matches the top-ranked IBM CUGA (60.3%), a production-level GPT-4.1–based agent [marreed2025towards], despite using the smaller open-source model DeepSeek-V3.1. With online adaptation, ReAct + ACE even surpasses IBM CUGA by 8.4% in TGC and 0.7% in SGC on the harder test-challenge split, underscoring the effectiveness of ACE in building comprehensive and self-evolving contexts for agents.
@@ 4.4 Results on Domain-Specific Benchmark
Table 2: Results on Financial Analysis Benchmark. "GT labels" indicates whether ground-truth labels are available to the Reflector during adaptation. With GT labels, ACE outperforms selected baselines by an average of 8.6%, highlighting the advantage of structured and evolving contexts for domain-specific reasoning. However, we also observe that in the absence of reliable feedback signals (e.g., ground-truth labels or execution outcomes), both ACE and other adaptive methods such as Dynamic Cheatsheet may degrade, suggesting that context adaptation depends critically on feedback quality.
Analysis.
@@
Moreover, we also observe that when ground-truth supervision or reliable execution signals are absent, both ACE and DC may degrade in performance. In such cases, the constructed context can be polluted by spurious or misleading signals, highlighting a potential limitation of inference-time adaptation without reliable feedback. This suggests that while ACE is robust under rich feedback (e.g., code execution results or formula correctness in agent tasks), its effectiveness depends on the availability of signals that allow the Reflector and Curator to make sound judgments. We return to this limitation in Appendix B.
4.5 Ablation Study
Table 3 reports ablation studies on the AppWorld benchmark, analyzing how individual design choices of ACE contribute to effective context adaptation. We examine three factors: (1) the Reflector with iterative refinement, our addition to the agentic framework beyond Dynamic Cheatsheet, (2) multi-epoch adaptation, which refines contexts over training samples multiple times, and (3) offline warmup, which initializes the context through offline adaptation before online adaptation begins.
@@ 4.6 Cost and Speed Analysis
Due to its support for incremental, “delta" context updates and non-LLM-based context merging and de-duplication, ACE demonstrates particular advantages in reducing the cost (in terms of the number of rollouts or the amount of dollar cost for token ingestion/generation) and latency of adaptation.
As examples, on the offline adaptation of AppWorld, ACE achieves 82.3% reduction in adaptation latency and 75.1% reduction in the number of rollouts as compared to GEPA (Table 4(a)). On the online adaptation of FiNER, ACE achieves 91.5% reduction in adaptation latency and 83.6% reduction in token dollar cost for token ingestion and generation as compared to DC (Table 4(b)).
@@ 5 Discussion
Longer Context $\neq$ Higher Serving Cost.
Although ACE produces longer contexts than methods such as GEPA, this does not translate to linearly higher inference cost or GPU memory usage. Modern serving infrastructures are increasingly optimized for long-context workloads through techniques such as the reuse [gim2024prompt, yao2025cacheblend], compression [liu2024kivi, liu2024cachegen], and offload [lee2024infinigen] of KV cache. These mechanisms allow frequently reused context segments to be cached locally or remotely, avoiding repetitive and expensive prefill operations. Ongoing advances in ML systems suggest that the amortized cost of handling long contexts will continue to decrease, making context-rich approaches like ACE increasingly practical in deployment.
Implications for Online and Continuous Learning.
@@
Online and continuous learning are key research directions in machine learning for addressing issues like distribution shifts [koh2021wilds, gulrajani2021domain] and limited training data [pan2010survey, hutchinson2017overcoming, zhuang2019transfer]. ACE offers a flexible and efficient alternative to conventional model fine-tuning, as adapting contexts is generally cheaper than updating model weights [brown2020gpt3, lester2021prompttuning, li2021prefixtuning, hu2021lora]. Moreover, because contexts are human-interpretable, ACE enables selective unlearning [cao2015unlearning, bourtoule2021sisa, liu2024rethinkingllmunlearning]—whether due to privacy or legal constraints [gdpr2016art17, ccpa1798_105], or when outdated or incorrect information is identified by domain experts. These are promising directions for future work, where ACE could play a central role in advancing continuous and responsible learning.
Appendix A Related Work on Agent Memory
A growing body of work explores how agents can accumulate experience from past trajectories and leverage external (often non-parametric) memory to guide future actions. AgentFly [zhou2025agentfly] presents an extensible framework where memory evolves continuously as agents solve tasks, enabling scalable reinforcement learning and long-horizon reasoning across diverse environments. AWM (Agent Workflow Memory) [wang2024agent] induces reusable workflows—structured routines distilled from past trajectories—and selectively injects them into memory to improve efficiency and generalization in web navigation benchmarks. A-MEM [xu2025mem] introduces a dynamically organized memory system inspired by the Zettelkasten method: each stored memory is annotated with structured attributes (e.g., tags, keywords, contextual descriptions) and automatically linked to relevant past entries, while existing entries are updated to integrate new knowledge, yielding adaptive and context-aware retrieval. Agentic Plan Caching [zhang2025cost] instead focuses on cost efficiency by extracting reusable plan templates from agent trajectories and caching them for fast execution at test time.
Together, these works demonstrate the value of external memory for improving adaptability, efficiency, and generalization in LLM agents. Our work differs by tackling the broader challenge of context adaptation, which spans not only agent memory but also system prompts, factual evidence, and other inputs underpinning AI systems. We further highlight two fundamental limitations of existing adaptation methods—brevity bias and context collapse—and show that addressing them is essential for robustness, reliability, and scalability beyond raw task performance. Accordingly, our evaluation considers not only accuracy but also cost, latency, and scalability.
Appendix B Limitations and Challenges

### v3

@@ 4.2 Baselines and Methods
Base LLM
The base model is evaluated directly on each benchmark without any context engineering, using the default prompts provided by dataset authors. For AppWorld, we follow the official ReAct (yao2023react) implementation released by the benchmark authors, and build all other baselines and methods on top of this framework.
In-Context Learning (ICL) (agarwal2024many)
@@
MIPROv2 (opsahl2024optimizing)
MIPROv2 is a popular prompt optimizer for LLM applications that works by jointly optimizing system instructions and in-context demonstrations via bayesian optimization. We use the official DSPy implementation (DSPyMIPROv2), setting auto="heavy" to maximize optimization performance.
GEPA (agrawal2025gepa)
GEPA (Genetic-Pareto) is a sample-efficient prompt optimizer based on reflective prompt evolution. It collects execution traces (reasoning, tool calls, intermediate outputs) and applies natural-language reflection to diagnose errors, assign credit, and propose prompt updates. A genetic Pareto search maintains a frontier of high-performing prompts, mitigating local optima. Empirically, GEPA outperforms reinforcement learning methods such as GRPO and prompt optimizers like MIPROv2, achieving up to 10–20% higher accuracy with as much as 35× fewer rollouts. We use the official DSPy implementation (DSPyGEPA), setting auto="heavy" to maximize optimization performance.
Dynamic Cheatsheet (DC) (suzgun2025dynamic)
DC is a test-time learning approach that introduces an adaptive external memory of reusable strategies and code snippets. By continuously updating this memory with newly encountered inputs and outputs, DC enables models to accumulate knowledge and reuse it across tasks, often leading to substantial improvements over static prompting methods. A key advantage of DC is that it does not require ground-truth labels: the model can curate its own memory from its generations, making the method highly flexible and broadly applicable. We use the official implementation released by the authors (Suzgun2025_DynamicCheatsheet_code) and set it to use the cumulative mode (DC-CU).
ACE (ours)
ACE optimizes LLM contexts for both offline and online adaptation through an agentic context engineering framework. To ensure fairness, we use the same LLM for the Generator, Reflector, and Curator (non-thinking mode of DeepSeek-V3.1 (deepseekai2024deepseekv3technicalreport)), preventing knowledge transfer from a stronger Reflector or Curator to a weaker Generator. This isolates the benefit of context construction itself. We additionally evaluate ACE with other backbone LLMs in the appendix, where we observe consistent gains. We adopt a batch size of 1 (constructing a delta context from each sample). We set the maximum number of Reflector refinement rounds and the maximum number of epoch in offline adaptation to 5.
4.3 Results on Agent Benchmark
@@
Table 1: Results on the AppWorld Agent Benchmark (DeepSeek-V3.1-671B as the Base LLM). “GT labels” indicates whether ground-truth labels are available to the Reflector during adaptation. We evaluate the ACE framework against multiple baselines on top of the official ReAct implementation, both for offline and online context adaptation. ReAct + ACE outperforms selected baselines by an average of 10.6%, and could achieve good performance even without access to GT labels.
Analysis: AppWorld
As shown in Table 1, ACE consistently improves over strong baselines on AppWorld. In the offline setting, ReAct + ACE outperforms both ReAct + ICL and ReAct + GEPA by significant margins (12.3% and 11.9%, respectively), demonstrating that structured, evolving, and detailed contexts enable more effective agent learning than fixed demonstrations or single optimized instruction prompts. These gains extend to the online setting, where ACE continues to outperform prior adaptive methods such as Dynamic Cheatsheet by an average of 7.6%.
@@
Notably, on the latest AppWorld leaderboard (as of September 20, 2025; Figure 5), ReAct + ACE (59.4% average) matches the top-1-ranked IBM CUGA (60.3%) [^1] We mention IBM CUGA as a rough contextual reference to show that ACE operates in a similar performance range on the AppWorld leaderboard. It is not used as a methodological baseline, and we do not make direct comparisons. CUGA’s internal design differs from ACE’s context-adaptation focus, and all baselines are evaluated under identical setups to isolate methodological effects rather than agent-engineering choices., a production-level GPT-4.1–based agent (marreed2025towards), despite using the much smaller open-source model DeepSeek-V3.1. With online adaptation, ReAct + ACE even surpasses IBM CUGA by 8.4% in TGC and 0.7% in SGC on test-challenge, underscoring the effectiveness of ACE in building comprehensive and self-evolving contexts for agents.
@@ 4.4 Results on Domain-Specific Benchmark
Table 2: Results on Financial Analysis Benchmark (DeepSeek-V3.1-671B as the Base LLM). “GT labels” indicates whether ground-truth labels are available to the Reflector during adaptation. With GT labels, ACE achieves consistent improvements in both offline and online settings, highlighting the advantage of structured and evolving contexts for domain-specific reasoning. However, we also observe that in the absence of reliable feedback signals (e.g., ground-truth labels or execution outcomes), both ACE and other adaptive methods such as Dynamic Cheatsheet may degrade, suggesting that context adaptation depends critically on feedback quality.
Analysis: Finance Benchmark
@@
Moreover, we also observe that when ground-truth supervision or reliable execution signals are absent, both ACE and DC may degrade in performance. In such cases, the constructed context can be polluted by spurious or misleading signals, highlighting a potential limitation of inference-time adaptation without reliable feedback. This suggests that while ACE is robust under rich feedback (e.g., code execution results or formula correctness in agent tasks), its effectiveness depends on the availability of signals that allow the Reflector and Curator to make sound judgments. We return to this limitation in §5.
Analysis: Medical and Text-to-SQL Benchmark
While this subsection focuses on finance as a detailed case study, ACE is not finance-specific: we also see consistent gains on other domain-specific tasks, including medical reasoning and text-to-SQL, suggesting that the same playbook-style context adaptation transfers across domains. Full results are reported in Appendix §A.2.
4.5 Generalization across LLMs
Table 1 and Table 2 use our default backbone (DeepSeek-V3.1), but ACE is not specific to this model. We can swap in other LLMs without changing the algorithm or prompts, and still see consistent gains on AppWorld and Finance benchmarks. Appendix §A.1 reports full results on GPT-OSS-120B, GPT-5.1, and Llama-3.3-70B-Instruct, where ACE improves over the corresponding base agents or models. These results suggest ACE is a generalizable method for test-time context evolution across LLM families.
4.6 Ablation Study and Sensitivity Analysis
Ablation Study
Table 3 reports ablation studies on AppWorld, analyzing how individual design choices of ACE contribute to effective context adaptation. We examine three factors: (1) the Reflector with iterative refinement, our addition to the agentic framework beyond Dynamic Cheatsheet, (2) multi-epoch adaptation, which refines contexts over training samples multiple times, and (3) offline warmup, which initializes the context through offline adaptation before online adaptation begins. Additionally, we study the effect of incremental context update and why it is a key enabler for ACE’s performance gain in Appendix §A.5.
@@ 4.6 Ablation Study and Sensitivity Analysis
Robustness to Reflection Quality
ACE is robust to reflection quality: it remains effective with a much weaker Reflector and shows only modest additional gains from stronger reflectors, and it degrades gracefully under noisy/harmful reflections, staying above the base model except under fully adversarial updates every iteration. Full experiment results are in Appendix §A.4.
Sensitivity to Hyperparameter Choice
ACE’s gains are stable across a wide range of reasonable hyperparameter settings (e.g., Reflector refinement rounds, number of adaptation epochs, and grow-and-refine thresholds): performance changes are modest, and ACE consistently remains above the corresponding baselines. Full discussion and detailed results are reported in Appendix §A.6.
@@ 4.7 Cost and Speed Analysis
Due to its support for incremental, “delta” context updates and non-LLM-based context merging and de-duplication, ACE demonstrates particular advantages in reducing the cost (in terms of the number of rollouts or the amount of dollar cost for token ingestion/generation) and latency of adaptation.
As examples, on the offline adaptation of AppWorld, ACE achieves 82.3% reduction in adaptation latency and 75.1% reduction in the number of rollouts as compared to GEPA (Table 4(a)). On the online adaptation of FiNER, ACE achieves 91.5% reduction in adaptation latency and 83.6% reduction in token dollar cost for token ingestion/generation as compared to DC (Table 4(b)).
@@
Fine-Grained Cost Analysis
We conduct a fine-grained cost analysis of ACE and GEPA (as a representative baseline). On AppWorld, ACE is substantially cheaper during offline adaptation, reducing input/output token usage by 80.8%/83.6% vs. GEPA: ACE avoids GEPA’s prompt-validation loop and replaces repeated full rewrites with localized delta updates. At evaluation time, while ACE may use more raw input tokens due to a richer playbook, this does not necessarily translate to higher billed serving cost because a large fraction of the context is reused by KV caching; we quantify this effect in the next paragraph. Full results, including a component-wise token breakdown for both methods, are in Appendix §A.3.
KV Cache Reuse: Longer Context $\neq$ Higher Serving Cost
Although ACE produces longer contexts than methods such as GEPA, this does not translate to linearly higher inference cost or GPU memory usage. Modern serving infrastructures are increasingly optimized for long-context workloads through techniques such as the reuse (gim2024prompt; yao2025cacheblend), compression (liu2024kivi; liu2024cachegen), and offload (lee2024infinigen; li2025continuum) of KV cache. These mechanisms allow frequently reused context segments to be cached locally or remotely, avoiding repetitive and expensive prefill operations. Ongoing advances in ML systems suggest that the amortized cost of handling long contexts is likely to decrease, making context-rich approaches like ACE increasingly practical in deployment. In our prompt-caching study with the OpenAI API (GPT-5.1), we find that ACE achieves high cache reuse: 91.8% of input tokens are served from cache during evaluation stage, which reduces billed input-token cost by 82.6% relative to counting raw context tokens.
@@ 5 Discussion
Implications for Online and Continual Learning
@@
Online and continual learning are key research directions in machine learning for addressing issues like distribution shifts (koh2021wilds; gulrajani2021domain) and limited training data (pan2010survey; hutchinson2017overcoming; zhuang2019transfer). ACE offers a flexible and efficient alternative to conventional model fine-tuning, as adapting contexts is generally cheaper than updating model weights (brown2020gpt3; lester2021prompttuning; li2021prefixtuning; hu2021lora). Moreover, because contexts are human-interpretable, ACE enables selective unlearning (cao2015unlearning; bourtoule2021sisa; liu2024rethinkingllmunlearning), whether due to privacy or legal constraints (gdpr2016art17; ccpa1798_105), or when outdated or incorrect information is identified by domain experts. These are promising directions for future work, where ACE could play a central role in advancing continuous and responsible learning.
Limitations and Challenges